# On the Performance of Dual-Antenna Repeater Assisted Bi-Static MIMO ISAC

### v1

@@ VI Numerical Results & Discussions
In this section, we quantify the effects of a repeater on the overall ISAC performance. The transmit power at the transmit BS is $1$ Watt. The path loss for the communication channels is modeled using the 33GPP Urban Microcell model. Each of the BSs are equipped with $8$ antennas. The height of the BS, repeaters, and users are $25$, $10$, and $1.5$ meters, respectively, following the models adopted in [Jianan_TWC]. The carrier frequency and bandwidth are set to $1.9$ GHz and $20$ MHz, respectively. At the receive BS, the noise spectral density is 174-174 dBm/Hz, and the noise figure is $9$ dB [Michail_TWC]. The repeater has the same noise power as the BS. The sensing-related parameters follow the Swerling-I model [Wei_ISAC_Channel_Model] and $\left\{\boldsymbol{a}_{{\mathsf{t}}},\boldsymbol{a}_{{\mathsf{r}}}\right\}$} are generated following [Behdad_Ozlem_TWC]. The clutter channels also follow the Microcell model; however, the channel gains are multiplied by an additional scaling to suppress the known parts of the target-free channels due to LOS and permanent obstacles [Behdad_Ozlem_TWC]. Other relevant parameters are also mentioned alongside the simulation results. The inter-BS channel and the residual interference is modeled according to [AC_CRM_TCOM_2024].
@@
In Fig. 3 illustrates the involved trade-offs between the different choices of precoders at the transmit BS. We can readily observe the degradation of the downlink SE for the target-centric precoder, as such a choice adds additional interference to the downlink users due to the dedicated target signals, for both with and without a repeater being present in the system. Communication-centric precoder nullifies that interference. Secondly, we also observe the benefits in the downlink SE due to the repeater.
@@ VII Remarks and Open Questions
This paper presented the initial results and a mathematical framework that facilitates repeater deployment for bi-static ISAC, and reported the benefits that can be obtained by using a repeater in the target hot-spot area. One interesting future study is to incorporate channel estimation within the developed framework. Furthermore, the impact of multiple repeaters (i.e. swarm-repeaters) in the system, including the effects of interaction between repeaters along with downlink power control, can be interesting to explore.
@@
$f(\alpha,\dot{{\mathbf{c}}};\mathcal{H}_{1})$
$=\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+{\mathbf{z}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}{\mathbf{z}}-2\Re{\left\{{\mathbf{z}}^{H}\mathbf{t}_{\mathcal{H}_{1}}\right\}},$
where ${\mathbf{z}}={\left[\alpha,\dot{{\mathbf{c}}}^{T}\right]}^{T}$ and the positive semi-definite (PSD) matrix ${\mathbf{Q}}_{\mathcal{H}_{1}}$ has the block-structure as shown in (16). Then, $\mathbf{t}_{\mathcal{H}_{1}}$ is expressed as shown in (12a). Now, it is easy to obtain an estimate of $\mathbf{z}$} from $f(\alpha,\dot{{\mathbf{c}}};\mathcal{H}_{1})$, which equals to ${\mathbf{Q}}_{\mathcal{H}_{1}}^{-1}\mathbf{t}_{\mathcal{H}_{1}}$, resulting

### v2

@@ VI Numerical Results & Discussions
In this section, we quantify the effects of a repeater on the overall ISAC performance. The transmit power at the transmit BS is $1$ Watt. The path loss for the communication channels is modeled using the 33GPP Urban Microcell model. Each of the BSs is equipped with $8$ antennas. The height of the BS, repeaters, and users are $25$, $10$, and $1.5$ meters, respectively, following the models adopted in [Jianan_TWC]. The carrier frequency and bandwidth are set to $1.9$ GHz and $20$ MHz, respectively. At the receive BS, the noise spectral density is 174-174 dBm/Hz, and the noise figure is $9$ dB [Michail_TWC]. The zone area is $500$ square meters, with the BSs being in $(0,250)$ and $(450,250)$; the sensing zone is $(225,250)$. The repeater is located at $(125,250)$. The user locations are randomly generated. The repeater has the same noise power as the BS. The sensing-related parameters follow the Swerling-I model [Wei_ISAC_Channel_Model] and $\left\{\boldsymbol{a}_{{\mathsf{t}}},\boldsymbol{a}_{{\mathsf{r}}}\right\}$} are generated following [Behdad_Ozlem_TWC]. For clutter, i.i.d. Gaussian taps are assumed. Other relevant parameters are also mentioned alongside the simulation results. The inter-BS channel and the residual interference is modeled according to [AC_CRM_TCOM_2024].
@@
Figure 4: Probability of detection with different repeater locations with respect to the target hotspot area. We assume the clutter variance to be $10$ dB below the noise floor and the PoFA is $0.01$.
@@
In Fig. 4 illustrates the involved trade-offs between the different choices of precoders at the transmit BS. The repeater gain is set to $5$ dB above the noise floor. We can readily observe the degradation of the downlink SE for the target-centric precoder, as such a choice adds additional interference to the downlink users due to the dedicated target signals, for both with and without a repeater being present in the system. Communication-centric precoder nullifies that interference. Secondly, we also observe the benefits in the downlink SE due to the repeater.
In Fig. 4, we illustrate the effects of moving the repeater towards the transmitter (varying the $x$-axis location) from the target hotspot area. This experiment reveals that a repeater near the hot-spot area can procure a high probability of detection value even for a lower target RCS. The reflected echo from a repeater, which is also further from the target, suffers severely from pathloss and contributes little to the target detection.
@@ VII Remarks and Open Questions
This paper presented the initial results and a mathematical framework that facilitates repeater deployment for bi-static ISAC, and reported the benefits that can be obtained by using a repeater in the target hot-spot area. One interesting future study is to incorporate channel estimation within the developed framework and include the time-varying nature of the channels. Furthermore, the impact of multiple repeaters (i.e. swarm-repeaters) in the system, including the effects of interaction between repeaters along with downlink power control, can be interesting to explore. A detailed comparison with other existing technologies, such as RIS, could also be examined as a future study.
@@
$\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+{\mathbf{z}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}{\mathbf{z}}-2\Re{\left\{{\mathbf{z}}^{H}\mathbf{t}_{\mathcal{H}_{1}}\right\}},$
where ${\mathbf{z}}={\left[\alpha,\dot{{\mathbf{c}}}^{T}\right]}^{T}$ and the positive semi-definite (PSD) matrix ${\mathbf{Q}}_{\mathcal{H}_{1}}$ has the block-structure as shown in (15). Then, $\mathbf{t}_{\mathcal{H}_{1}}$ is expressed as shown in (11a). Now, the estimate of $\mathbf{z}$} obtained from $f(\alpha,\dot{{\mathbf{c}}};\mathcal{H}_{1})$ equals to ${\mathbf{Q}}_{\mathcal{H}_{1}}^{-1}\mathbf{t}_{\mathcal{H}_{1}}$, resulting